# Evidence of Energy Injection in the Short and Distant GRB 250221A

### v4

@@ 6 Summary & Discussion
The light curve shows a typical decay in the optical and X-ray frequencies before $T<10^{4}$ s for synchrotron emission, with the characteristic frequencies being $\nu_{m}\lesssim\nu_{\rm opt}\lesssim\nu_{X}\lesssim\nu_{c}$ indicating a slow-cooling synchrotron regime in a circumstellar profile of constant density. At later times, after $T\sim 0.6$ d, we observe an excess in the X-ray, optical, and radio light curves.
@@
Regardless, the central engine can be argued to be either a millisecond magnetar (Metzger et al., 2011) or a hyperaccreting black hole (BH) (MacFadyen & Woosley, 1999b) that should be able to provide a total energy $E_{\rm tot}=E_{k}+E_{\gamma}\approx E_{k}=4.5\times 10^{51}$ erg to power the GRB and its afterglow. In the Collapsar scenario, the total duration over which the outflow is powered is obtained from $t_{\rm jet}=t_{\rm bo}+T_{90}/(1+z)\sim 10$ s where we take $t_{\rm bo}\sim 9$ s for the jet to break out of the stellar progenitor (Bromberg & Tchekhovskoy, 2016). The mean jet power that the engine must be able to provide is $L_{\rm jet}=E_{\rm tot}/\eta_{j}t_{\rm jet}\sim 4.5\times 10^{51}\eta_{j,-1}^{-1}t_{\rm jet,1}^{-1}\,{\rm erg\,s}^{-1}$ for an assumed efficiency of $\eta_{j}=0.1\eta_{\rm j,-1}$. In the case of a binary merger of two NSs, the jet breakout time out of the dynamical ejecta is $t_{\rm bo}\sim 0.5$ s (e.g., Moharana & Piran, 2017; Gill et al., 2019), which would demand a mean jet power $L_{\rm jet}\sim 3\times 10^{52}\eta_{j,-1}^{-1}(t_{\rm jet}/1.5\,{\rm s})^{-1}\,{\rm erg\,s}^{-1}$. The total energy and mean jet power in this event can be supplied by any of the various known jet launching mechanisms, namely $\nu\bar{\nu}$ annihilation from accretion onto a NS or BH (see, e.g. MacFadyen & Woosley, 1999b; Zalamea & Beloborodov, 2011; Globus & Levinson, 2014), an MHD wind from a millisecond magnetar (Metzger et al., 2011), a Kerr BH with a Blandford-Znajek engine (Blandford & Znajek, 1977; Brown et al., 2000), or a Schwarzschild/Kerr BH with a BP engine (Blandford & Payne, 1982).
@@
We attribute the rebrightening in the afterglow light curve at $T>0.6$ days to energy injection by the central engine into the forward shock produced by a relativistic jet. Three popular scenarios of energy injection in which a single and isolated rebrightening episode is realized are refreshing of the forward shock due to a mild collision (as opposed to a violent one) between two matter shells (Kumar & Piran, 2000), the ejecta having a radial velocity stratification (Rees & Mészáros, 1998; Sari & Mészáros, 2000), or energy injected by a spinning down millisecond magnetar via an MHD wind (Zhang & Mészáros, 2001). In all of these three scenarios, the energy of the blast wave is gently increased without the formation of strong shocks propagating through both the initial blast wave and energizing material. The main result is an increase in the emission from the forward shock produced by the initial blast wave. Alternatively, a millisecond-magnetar central engine can power the initial jet, and energy can be later injected into it by the gravitational collapse into a Schwarzschild BH as the magnetar spins down and loses centrifugal support, thus releasing the remaining binding energy $E_{B}\simeq 5\times 10^{53}{\rm erg}$ (see Moreno Méndez et al., 2015, for further discussion and details on this model; see also Appendix B).
@@
The complete multi-wavelength photometric follow-up and modelling of GRB 250221A provide key insights into its progenitor, central engine, and energy injection mechanisms, while afterglow spectroscopy proved crucial for characterizing the CBM of this short-duration burst. Together, these observations add to the growing number of long and short GRBs that exhibit rebrightening episodes in their afterglows, commonly attributed to refreshed shocks and late-time central engine activity (e.g. Granot et al., 2003; Fan & Xu, 2006; Soderberg et al., 2006; de Ugarte Postigo et al., 2007; Troja et al., 2007; Rowlinson et al., 2013; Hascoët et al., 2012; Laskar et al., 2015, 2018; Lamb et al., 2019; de Wet et al., 2023; Moss et al., 2023; de Wet et al., 2024; Schroeder et al., 2024, 2025). The case of GRB 250221A demonstrates how coordinated, multi-wavelength follow-up and prompt spectroscopy can reveal new facets of GRB physics and challenge our current understanding of their progenitors.
@@ Appendix A Refreshed Shock from Two Shell Collision
$t_{\rm em}\approx\left[1-\frac{(1+m)\Gamma_{3}^{2}}{\Gamma_{0,6}^{2}}\right]T_{\rm bright}\,$, (15)
@@
$\left(\frac{\Gamma_{0,3}}{\Gamma_{0,6}}\right)^{2}=\left[1-(1+z)\frac{t_{\rm em}}{T_{\rm bright}}\right]\frac{[(1+m)\hat{T}_{\rm bright}]^{m\over 1+m}}{(1+m)}\,$, (16)
@@
$\left(\frac{\Gamma_{0,3}}{\Gamma_{0,6}}\right)^{2}=\frac{[(1+m)\hat{T}_{\rm bright}]^{m\over 1+m}}{(1+m)}\,.$ (17)
@@
Since $t_{\rm em}/T_{\rm bright}<T_{90}/T_{\rm bright}\ll 1$ is guaranteed in a binary merger scenario, it forces $\Gamma_{3}/\Gamma_{0,6}\approx(1+m)^{-1/2}=1/2$ for $m=3$ (Kumar & Piran, 2000). This also implies that the delay between the emission time of the shells can be approximated as negligible, with $t_{\rm em}\approx 0$. The arrival time of radiation from the collision radius is given by $\hat{T}_{\rm bright}=\hat{R}_{\rm coll}^{1+m}/(1+m)$ when $\Gamma_{3}\gg 1$, where
$\hat{R}_{\rm coll}\equiv\frac{R_{\rm coll}}{R_{\rm dec}}=\left[(1+m)\left(\frac{\Gamma_{0,3}}{\Gamma_{0,6}}\right)^{2}\right]^{1/m}\,.$ (18)
In this case it is required that $\Gamma_{0,6}<\Gamma_{0,3}$ for $\hat{R}_{\rm coll}>1$.
@@
The relative LF between the two shells at the time of collision is $\Gamma_{\rm rel}\simeq(\Gamma_{3}/\Gamma_{0,6}+\Gamma_{0,6}/\Gamma_{3})/2=5/4$ for $\Gamma_{3},\,\Gamma_{0,6}\gg 1$. In order to produce a strong shock, $\Gamma_{\rm rel}$ must significantly exceed the LF corresponding to the local sound speed ($c_{s}$) in the first shell. Since the collision occurs after the passage of the reverse shock in the first shell, due to its interaction with the external medium, it can be assumed to be relativistically hot. In that case, the sound speed is $\beta_{s}=c_{s}/c=1/\sqrt{3}$, with the corresponding LF $\Gamma_{s}=\sqrt{3/2}\simeq 1.22\lesssim\Gamma_{\rm rel}$.
Figure 12: Sketch (not to scale) of the different shocked and unshocked regions that arise in the collision of two kinetic-energy-dominated shells. The thick solid line indicates the pressure, or the internal energy density, in each region. The thin solid line indicates the comoving mass density. Figure adapted from Zhang & Mészáros (2002).
@@
Figure 13: Shell dynamics with $E_{\rm k,iso,6}/E_{\rm k,iso,3}=24$ and $\Gamma_{0,3}/\Gamma_{0,6}=46.5$. All other parameters are the same as the best-fit parameters shown in the legend of Figure 8. Different important radii are also shown, where the first shell decelerates at $R_{\rm dec}$ and the second shell catches up from behind and merges with the first at $R_{\rm coll}$. Two shocks, in addition to the external forward shock by the first shell, form at $R_{\rm sh}$, from which a reverse shock crosses the second shell at $R_{\Delta}$. The dynamics of the first decelerating shell are altered (dashed line) at $R>R_{\Delta}$ due to injection of energy by the second shell.
@@
Next, we calculate the emission that arises from the two shell collision and which produces the re-brightening. As the two shells collide, a forward shock is launched into the first shell and a reverse shock into the second shell. In total, there are six different regions with different dynamical evolution as illustrated in Figure 12 (Zhang & Mészáros, 2002): (1) unshocked ISM, (2) ISM shocked by the forward shock from the first shell, (3) material from the first shell shocked by the early passage ($T_{\times 1}\lesssim T_{\rm dec}$) of the reverse shock, (4) relativistically hot material in the first shell now shocked by the forward shock from the second shell, (5) material of the second shell now shocked by the reverse shock from the collision, and (6) unshocked material of the second shell. All physical quantities in what follows refer to these regions with the corresponding number. To relate the properties of the upstream flow with that of the downstream across the three shocks, we use the shock jump conditions. For a strong shock going into a cold medium, i.e. $\rho_{u}c^{2}\gg e_{u}+P_{u}$, where $\rho_{u}$, $e_{u}$, and $P_{u}$ are the proper mass density, internal energy density, and pressure of the upstream medium (with subscript ‘u’), the shock-jump conditions yield (Blandford & McKee, 1976)
@@
where $\Gamma_{ud}$ is the relative LF between the upstream and downstream (with subscript ‘d’) media and $\hat{\gamma}$ is the adiabatic index of the downstream medium with $\hat{\gamma}=4/3\,(5/3)$ for a relativistic (non-relativistic) fluid. The above shock-jump condition applies to the forward shock between regions (1) and (2) and the reverse shock between regions (5) and (6). The shock between regions (3) and (4) is propagating into a relativistically hot region, i.e. $e_{u}+P_{u}\gg\rho_{u}c^{2}$ and $P_{u}=(\hat{\gamma}-1)e_{u}=e_{u}/3$ for $\hat{\gamma}=4/3$. Therefore, it admits different shock-jump conditions (Kumar & Piran, 2000; Zhang & Mészáros, 2002)
@@
Since the emission from the reverse shock makes the dominant contribution to the total flux until the shock crosses the second shell, we first derive the flux scaling relations for that. When the collision occurs and the second shell decelerates, it is important to know if the reverse shock is propagating into a radially expanding shell (thin shell regime) or that with a fixed radial width (thick shell regime; Sari & Piran 1995). If the initial radial width of the second shell is $\Delta_{0,6}=c\delta T/(1+z)$, where $\delta T$ is the activity time of the central engine, the shell starts to spread at the radius $R_{s}=\Delta_{0,6}^{\prime}\Gamma_{0,6}=\Delta_{0,6}\Gamma_{0,6}^{2}$. For the shell to be thick at the time of collision, i.e. $\hat{R}_{s}>\hat{R}_{\rm coll}$, its initial LF must exceed that of the first shell,
$\frac{\Gamma_{0,6}}{\Gamma_{0,3}}>\sqrt{\frac{2T_{\rm dec}}{\delta T}}[(1+m)\hat{T}_{\rm bright}]^{1\over 2(1+m)}\simeq 16.8\,T_{\rm bright,5}^{1/8}\,T_{\rm dec,1}^{3/8}\,\delta T^{-1/2}\,$, (21)
@@
which contradicts the earlier condition of $\Gamma_{0,2}<\Gamma_{0,1}$. Therefore, the second shell must start to spread before the collision, in which case the comoving number density of the shell is $n_{6}=M_{6}\Gamma_{0,6}/4\pi R^{3}m_{p}\propto R^{-3}$, where $M_{6}=E_{\rm k,iso,6}/\Gamma_{0,6}c^{2}$ is the shell mass and $E_{\rm k,iso,6}$ is its kinetic energy.
@@
From the shock-jump condition in Equation (19), we find $n_{5}=4\Gamma_{56}n_{6}$ and $e_{5}=4(\Gamma_{56}-1)\Gamma_{56}n_{6}m_{p}c^{2}$. Regions 4 and 5 are separated by a contact discontinuity, where the shocked fluids in the two regions are in pressure equilibrium, i.e. $P_{5}=P_{4}$, and also move with the same velocity, i.e. $\Gamma_{5}=\Gamma_{4}$, which implies that $\Gamma_{56}=\Gamma_{46}$. In order to calculate $\Gamma_{56}$ we need to know $\Gamma_{4}$, where the latter is obtained from the shock-jump condition in Equation (20) that yields, $\Gamma_{43}^{2}=(3e_{4}/e_{3}+1)(e_{4}/e_{3}+3)/16(e_{4}/e_{3})$. Here, $e_{4}=e_{5}$ due to both being relativistically hot and in pressure equilibrium. We assume that region 3 is still relativistically hot after RS passage, so that $P_{3}\approx P_{2}$ and $e_{3}\approx e_{2}$, and also $\Gamma_{3}\approx\Gamma_{2}$. In that case, the jump condition in Equation (19) yields $e_{4}/e_{3}\approx e_{5}/e_{2}\approx(\Gamma_{56}-1)\Gamma_{56}n_{6}/\Gamma_{2}^{2}n_{1}$.
@@
We solve the coupled dynamics of the different regions and show the LFs of the relevant regions in the left column of Figure 13 for the two different solutions obtained from the light curve fits in Figure 8. Even though the two shells make contact at the radius $R_{\rm coll}$, the shocks do not form until the shells reach $R_{\rm shock}\geq R_{\rm coll}$ when the condition that $e_{4}>e_{3}$ is met. At $R>R_{\rm shock}$, the RS starts to propagate into the inner shell and extracts its kinetic energy. The differential number of baryons entering the shock can be calculated as $dN_{6}=4\pi R^{2}n_{6}\Gamma_{6}d\Delta_{5}$, where $d\Delta_{5}=(\beta_{6}-\beta_{5})(1-\Gamma_{6}n_{6}/\Gamma_{5}n_{5})^{-1}dR$ is the infinitesimal width of the shocked region, as obtained from baryon number conservation. The total mass shocked by the RS is then $M_{5}=\int_{R_{\rm shock}}^{R_{\Delta}}m_{p}dN_{6}$, where $R_{\Delta}$ is the shock crossing radius, and it is obtained when $M_{5}=M_{6}$. In the cases shown in Figure 13, the RS moves through the inner shell at a non-relativistic speed, with $\Gamma_{\rm RS}\sim\Gamma_{56}$, and crosses the shell over $\Delta R/R_{\rm cross}\lesssim 1.5-2$. After RS passage, it is expected that the LF of the now entirely shocked inner shell follows $\Gamma_{5}\propto R^{-g}$, where $3/2\leq g\leq 7/2$ (Kobayashi & Sari, 2000) and we assume $g=2$. Due to energy injection by the inner shell, the dynamics of the merged two-shell system are altered, which requires more careful modelling and are not considered here.

### v5

@@ 6 Summary & Discussion
The light curve shows a typical decay in the optical and X-ray frequencies before $T<3\times 10^{4}$ s for synchrotron emission, with the characteristic frequencies being $\nu_{m}\lesssim\nu_{\rm opt}\lesssim\nu_{X}\lesssim\nu_{c}$ indicating a slow-cooling synchrotron regime in a circumstellar profile of constant density. At later times, after $T\sim 0.6$ d, we observe an excess in the X-ray, optical, and radio light curves.
@@
Regardless, the central engine can be argued to be either a millisecond magnetar (Metzger et al., 2011) or a hyperaccreting black hole (BH) (MacFadyen and Woosley, 1999a) that should be able to provide a total energy $E_{\rm tot}=E_{k}+E_{\gamma}\approx E_{k}\simeq 4.5\times 10^{51}$ erg to power the GRB and its afterglow. In the Collapsar scenario, the total duration over which the outflow is powered is obtained from $t_{\rm jet}=t_{\rm bo}+T_{90}/(1+z)\sim 10$ s where we take $t_{\rm bo}\sim 9$ s for the jet to break out of the stellar progenitor (Bromberg and Tchekhovskoy, 2016). The mean jet power that the engine must be able to provide is $L_{\rm jet}=E_{\rm tot}/\eta_{j}t_{\rm jet}\sim 4.5\times 10^{51}\eta_{j,-1}^{-1}t_{\rm jet,1}^{-1}\,{\rm erg\,s}^{-1}$ for an assumed efficiency of $\eta_{j}=0.1\eta_{\rm j,-1}$. In the case of a binary merger of two NSs, the jet breakout time out of the dynamical ejecta is $t_{\rm bo}\sim 0.5$ s (e.g., Moharana and Piran, 2017; Gill et al., 2019), which would demand a mean jet power $L_{\rm jet}\sim 3\times 10^{52}\eta_{j,-1}^{-1}(t_{\rm jet}/1.5\,{\rm s})^{-1}\,{\rm erg\,s}^{-1}$. The total energy and mean jet power in this event can be supplied by any of the various known jet launching mechanisms, namely $\nu\bar{\nu}$ annihilation from accretion onto a NS or BH (see, e.g. MacFadyen and Woosley, 1999a; Zalamea and Beloborodov, 2011; Globus and Levinson, 2014), an MHD wind from a millisecond magnetar (Metzger et al., 2011), a Kerr BH with a Blandford-Znajek engine (Blandford and Znajek, 1977; Brown et al., 2000), or a Schwarzschild/Kerr BH with a Blandford–Payne engine (Blandford and Payne, 1982).
@@
We attribute the rebrightening in the afterglow light curve at $T>0.6$ days to energy injection by the central engine into the forward shock produced by a relativistic jet. Three popular scenarios of energy injection in which a single and isolated rebrightening episode is realized are refreshing of the forward shock due to a mild collision (as opposed to a violent one) between two matter shells (Kumar and Piran, 2000), the ejecta having a radial velocity stratification (Rees and Mészáros, 1998; Sari and Mészáros, 2000), or energy injected by a spinning down millisecond magnetar via an MHD wind (Zhang and Mészáros, 2001). In all of these three scenarios, the energy of the blast wave is gently increased without the formation of strong shocks propagating through both the initial blast wave and energizing material. The main result is an increase in the emission from the forward shock produced by the initial blast wave. Alternatively, the merger of two NSs produces a hypermassive (beyond the TOV limit) millisecond-magnetar central engine which can power the initial jet. Energy can be later injected into the GRB jet by the gravitational collapse into a Schwarzschild BH, as the magnetar spins down and loses centrifugal support, thus releasing the remaining binding energy $E_{B}\simeq 5\times 10^{53}{\rm erg}$ (see Moreno Méndez et al., 2015, for further discussion and details on this model; see also Appendix B) into a second shock, channelled into the pathway left by the initial jet, that eventually collides with the slowed-down head of the jet.
Rebrightenings of this kind are extremely rare among short GRBs. Although some short bursts show early-time X-ray or optical plateaus—often interpreted as signatures of long-lived central engines—clear late-time rises in the afterglow flux are not typically observed. A few events, such as GRB 060614 (Yang et al., 2015) and GRB 231117A (Schroeder et al., 2025b), exhibit shallow decay phases lasting up to $\sim$1 day, but these do not display the distinct, pronounced rebrightening seen in GRB 250221A; instead, their afterglows remain nearly constant or transition smoothly into the standard decay regime. In addition to this unusual behaviour, GRB 250221A ranks among the most luminous and energetic short-GRB afterglows at $T=1$ day, further distinguishing it from the known population. This behaviour contrasts sharply with that of several long GRBs, where rises on similar timescales have been linked to supernova emission, refreshed shocks, or structured jets. If GRB 250221A indeed belongs to the Type I class according to its prompt properties, the late-time rise we observe would represent the first clear instance of such behaviour within the short-GRB population (see Fig. 10 from Kann et al., 2011), underscoring the unusual nature of this event and suggesting that an atypical energy-injection or density-structure scenario may be required.
The complete multi-wavelength photometric follow-up and modelling of GRB 250221A provide key insights into its progenitor, central engine, and energy injection mechanisms, while afterglow spectroscopy proved crucial for the identification of the host galaxy (and therefore properties such as the SFR and distance).
@@
Together, these observations add to the growing number of long and short GRBs that exhibit rebrightening episodes in their afterglows, commonly attributed to refreshed shocks and late-time central engine activity (e.g. Granot et al., 2003; Fan and Xu, 2006; Soderberg et al., 2006; de Ugarte Postigo et al., 2007; Troja et al., 2007; Rowlinson et al., 2013; Hascoët et al., 2012; Laskar et al., 2015, 2018; Lamb et al., 2019; de Wet et al., 2023; Moss et al., 2023; de Wet et al., 2024; Schroeder et al., 2024, 2025a). The case of GRB 250221A demonstrates how coordinated, multi-wavelength follow-up and prompt afterglow spectroscopy can reveal new facets of GRB physics and challenge our current understanding of their progenitors.
@@ Acknowledgements
We are grateful to the referee for their thorough reading of the manuscript and for their insightful and constructive comments, which helped improve the presentation and robustness of our results.
@@ Appendix A Refreshed Shock from Two Shell Collision
$t_{\rm em}\approx\left[1-\frac{(1+m)\Gamma_{(3)}^{2}}{\Gamma_{(6),0}^{2}}\right]T_{\rm bright}\,$, (15)
@@
$\left(\frac{\Gamma_{(3),0}}{\Gamma_{(6),0}}\right)^{2}=\left[1-(1+z)\frac{t_{\rm em}}{T_{\rm bright}}\right]\frac{[(1+m)\hat{T}_{\rm bright}]^{m\over 1+m}}{(1+m)}\,$, (16)
@@
$\left(\frac{\Gamma_{(3),0}}{\Gamma_{(6),0}}\right)^{2}=\frac{[(1+m)\hat{T}_{\rm bright}]^{m\over 1+m}}{(1+m)}\,.$ (17)
@@
Since $t_{\rm em}/T_{\rm bright}<T_{90}/T_{\rm bright}\ll 1$ is guaranteed in a binary merger scenario, it forces $\Gamma_{(3)}/\Gamma_{(6),0}\approx(1+m)^{-1/2}=1/2$ for $m=3$ (Kumar and Piran, 2000). This also implies that the delay between the emission time of the shells can be approximated as negligible, with $t_{\rm em}\approx 0$. The arrival time of radiation from the collision radius is given by $\hat{T}_{\rm bright}=\hat{R}_{\rm coll}^{1+m}/(1+m)$ when $\Gamma_{(3)}\gg 1$, where
$\hat{R}_{\rm coll}\equiv\frac{R_{\rm coll}}{R_{\rm dec}}=\left[(1+m)\left(\frac{\Gamma_{(3),0}}{\Gamma_{(6),0}}\right)^{2}\right]^{1/m}\,.$ (18)
In this case it is required that $\Gamma_{(6),0}<\Gamma_{(3),0}$ for $\hat{R}_{\rm coll}>1$.
@@
The relative LF between the two shells at the time of collision is $\Gamma_{\rm rel}\simeq(\Gamma_{(3)}/\Gamma_{(6),0}+\Gamma_{(6),0}/\Gamma_{(3)})/2=5/4$ for $\Gamma_{(3)},\,\Gamma_{(6),0}\gg 1$. In order to produce a strong shock, $\Gamma_{\rm rel}$ must significantly exceed the LF corresponding to the local sound speed ($c_{s}$) in the first shell. Since the collision occurs after the passage of the reverse shock in the first shell, due to its interaction with the external medium, it can be assumed to be relativistically hot. In that case, the sound speed is $\beta_{s}=c_{s}/c=1/\sqrt{3}$, with the corresponding LF $\Gamma_{s}=\sqrt{3/2}\simeq 1.22\lesssim\Gamma_{\rm rel}$.
@@
Figure 13: Shell dynamics with $E_{(6),\rm k,iso}/E_{(3),\rm k,iso}=24$ and $\Gamma_{(3),0}/\Gamma_{(6),0}=46.5$. All other parameters are the same as the best-fit parameters shown in the legend of Figure 8. Different important radii are also shown, where the first shell decelerates at $R_{\rm dec}$ and the second shell catches up from behind and merges with the first at $R_{\rm coll}$. Two shocks, in addition to the external forward shock by the first shell, form at $R_{\rm shock}$, from which a reverse shock crosses the second shell at $R_{\Delta}$. The dynamics of the first decelerating shell are altered (dashed line) at $R>R_{\Delta}$ due to injection of energy by the second shell.
@@
Next, we calculate the emission that arises from the two shell collision and which produces the re-brightening. As the two shells collide, a forward shock is launched into the first shell and a reverse shock into the second shell. In total, there are six different regions with different dynamical evolution as illustrated in Figure 10 (Zhang and Mészáros, 2002): (1) unshocked ISM, (2) ISM shocked by the forward shock from the first shell, (3) material from the first shell shocked by the early passage ($T_{\times 1}\lesssim T_{\rm dec}$) of the reverse shock, (4) relativistically hot material in the first shell now shocked by the forward shock from the second shell, (5) material of the second shell now shocked by the reverse shock from the collision, and (6) unshocked material of the second shell. All physical quantities in what follows refer to these regions with the corresponding number. To relate the properties of the upstream flow with that of the downstream across the three shocks, we use the shock jump conditions. For a strong shock going into a cold medium, i.e. $\rho_{u}c^{2}\gg e_{u}+P_{u}$, where $\rho_{u}$, $e_{u}$, and $P_{u}$ are the proper mass density, internal energy density, and pressure of the upstream medium (with subscript ‘u’), the shock-jump conditions yield (Blandford and McKee, 1976)
@@
where $\Gamma_{ud}$ is the relative LF between the upstream and downstream (with subscript ‘d’) media and $\hat{\gamma}$ is the adiabatic index of the downstream medium with $\hat{\gamma}=4/3\,(5/3)$ for a relativistic (non-relativistic) fluid. The above shock-jump condition applies to the forward shock between regions (1) and (2) and the reverse shock between regions (5) and (6). The shock between regions (3) and (4) is propagating into a relativistically hot region, i.e. $e_{u}+P_{u}\gg\rho_{u}c^{2}$ and $P_{u}=(\hat{\gamma}-1)e_{u}=e_{u}/3$ for $\hat{\gamma}=4/3$. Therefore, it admits different shock-jump conditions (Kumar and Piran, 2000; Zhang and Mészáros, 2002)
@@
Since the emission from the reverse shock makes the dominant contribution to the total flux until the shock crosses the second shell, we first derive the flux scaling relations for that. When the collision occurs and the second shell decelerates, it is important to know if the reverse shock is propagating into a radially expanding shell (thin shell regime) or that with a fixed radial width (thick shell regime; Sari and Piran 1995). If the initial radial width of the second shell is $\Delta_{(6),0}=c\delta T/(1+z)$, where $\delta T$ is the activity time of the central engine, the shell starts to spread at the radius $R_{s}=\Delta_{(6),0}^{\prime}\Gamma_{(6),0}=\Delta_{(6),0}\Gamma_{(6),0}^{2}$. For the shell to be thick at the time of collision, i.e. $\hat{R}_{s}>\hat{R}_{\rm coll}$, its initial LF must exceed that of the first shell,
$\frac{\Gamma_{(6),0}}{\Gamma_{(3),0}}$ $>\sqrt{\frac{2T_{\rm dec}}{\delta T}}[(1+m)\hat{T}_{\rm bright}]^{1\over 2(1+m)}$
$\simeq 16.8\,T_{\rm bright,5}^{1/8}\,T_{\rm dec,1}^{3/8}\,\delta T^{-1/2}\,$,
@@
which contradicts the earlier condition of $\Gamma_{(2),0}<\Gamma_{(1),0}$. Therefore, the second shell must start to spread before the collision, in which case the comoving number density of the shell is $n_{(6)}=M_{(6)}\Gamma_{(6),0}/4\pi R^{3}m_{p}\propto R^{-3}$, where $M_{(6)}=E_{(6),\rm k,iso}/\Gamma_{(6),0}c^{2}$ is the shell mass and $E_{(6),\rm k,iso}$ is its kinetic energy.
@@
From the shock-jump condition in Equation (19), we find $n_{(5)}=4\Gamma_{(5)(6)}n_{(6)}$ and $e_{(5)}=4(\Gamma_{(5)(6)}-1)\Gamma_{(5)(6)}n_{(6)}m_{p}c^{2}$. Regions (4) and (5) are separated by a contact discontinuity, where the shocked fluids in the two regions are in pressure equilibrium, i.e. $P_{(5)}=P_{(4)}$, and also move with the same velocity, i.e. $\Gamma_{(5)}=\Gamma_{(4)}$, which implies that $\Gamma_{(5)(6)}=\Gamma_{(4)(6)}$. In order to calculate $\Gamma_{(5)(6)}$ we need to know $\Gamma_{(4)}$, where the latter is obtained from the shock-jump condition in Equation (20) that yields, $\Gamma_{(4)(3)}^{2}=(3e_{(4)}/e_{(3)}+1)(e_{(4)}/e_{(3)}+3)/16(e_{(4)}/e_{(3)})$. Here, $e_{(4)}=e_{(5)}$ due to both being relativistically hot and in pressure equilibrium. We assume that region (3) is still relativistically hot after RS passage, so that $P_{(3)}\approx P_{(2)}$ and $e_{(3)}\approx e_{(2)}$, and also $\Gamma_{(3)}\approx\Gamma_{(2)}$. In that case, the jump condition in Equation (19) yields $e_{(4)}/e_{(3)}\approx e_{(5)}/e_{(2)}\approx(\Gamma_{(5)(6)}-1)\Gamma_{(5)(6)}n_{(6)}/\Gamma_{(2)}^{2}n_{(1)}$.
@@
We solve the coupled dynamics of the different regions and show the LFs of the relevant regions in the left column of Figure 13 for the two different solutions obtained from the light curve fits in Figure 8. Even though the two shells make contact at the radius $R_{\rm coll}$, the shocks do not form until the shells reach $R_{\rm shock}\geq R_{\rm coll}$ when the condition that $e_{(4)}>e_{(3)}$ is met. At $R>R_{\rm shock}$, the RS starts to propagate into the inner shell and extracts its kinetic energy. The differential number of baryons entering the shock can be calculated as $dN_{(6)}=4\pi R^{2}n_{(6)}\Gamma_{(6)}d\Delta_{(5)}$, where $d\Delta_{(5)}=(\beta_{(6)}-\beta_{(5)})(1-\Gamma_{(6)}n_{(6)}/\Gamma_{(5)}n_{(5)})^{-1}dR$ is the infinitesimal width of the shocked region, as obtained from baryon number conservation. The total mass shocked by the RS is then $M_{(5)}=\int_{R_{\rm shock}}^{R_{\Delta}}m_{p}dN_{(6)}$, where $R_{\Delta}$ is the shock crossing radius, and it is obtained when $M_{(5)}=M_{(6)}$. In the cases shown in Figure 13, the RS moves through the inner shell at a non-relativistic speed, with $\Gamma_{\rm RS}\sim\Gamma_{(5)(6)}$, and crosses the shell over $\Delta R/R_{\rm cross}\lesssim 1.5-2$. After RS passage, it is expected that the LF of the now entirely shocked inner shell follows $\Gamma_{(5)}\propto R^{-g}$, where $3/2\leq g\leq 7/2$ (Kobayashi and Sari, 2000) and we assume $g=2$. Due to energy injection by the inner shell, the dynamics of the merged two-shell system are altered, which requires more careful modelling and are not considered here.